Income Statements under Absorption Costing and Variable Costing

Joplin Industries Inc. manufactures and sells high-quality sporting goods equipment under its highly recognizable J-Sports logo. The company began operations on May 1 and operated at 100% of capacity (39,600 units) during the first month, creating an ending inventory of 3,600 units. During June, the company produced 36,000 garments during the month but sold 39,600 units at $115 per unit. The June manufacturing costs and selling and administrative expenses were as follows:

Number of Units Unit Cost Total
Cost
Manufacturing costs in June 1 beginning inventory:
Variable 3,600 $46.00 $165,600
Fixed 3,600 17.00 61,200
Total $63.00 $226,800
Manufacturing costs in June:
Variable 36,000 $46.00 $1,656,000
Fixed 36,000 18.70 673,200
Total $64.70 $2,329,200
Selling and administrative expenses in June:
Variable 39,600 22.10 $875,160
Fixed 39,600 7.00 277,200
Total 29.10 $1,152,360

a. Prepare an income statement according to the absorption costing concept for June.

Joplin Industries Inc.
Absorption Costing Income Statement
For the Month Ended June 30
$
Cost of goods sold:
$
$
$

b. Prepare an income statement according to the variable costing concept for June.

Joplin Industries Inc.
Variable Costing Income Statement
For the Month Ended June 30
$
$
$
Fixed costs:
$
$

c. What is the reason for the difference in the amount of income from operations reported in (a) and (b)?

Under the   method, the fixed manufacturing cost included in the cost of goods sold is matched with the revenues. Under  , all of the fixed manufacturing cost is deducted in the period in which it is incurred, regardless of the amount of inventory change. Thus, when inventory decreases, the   income statement will have a lower income from operations.
Answer #1

a.

Sales Revenue $       45,54,000 =39600*115
Cost of Goods Sold $       25,56,000 =226800+2329200
Gross Profit $       19,98,000
Selling and administrative expenses $       11,52,360
Operating Income $         8,45,640

b.

Sales Revenue $       45,54,000 =39600*115
Less : Variable Costs
Cost of Goods Sold $       18,21,600 =165600+1656000
Selling and administrative expenses $         8,75,160
Total Variable Costs $       26,96,760
Contribution Margin $       18,57,240
Less : Fixed Costs
Cost of Goods Sold $         6,73,200
Selling and administrative expenses $         2,77,200
Total Fixed Costs $         9,50,400
Operating Income $         9,06,840

c.

Under the absorption costing method, the fixed manufacturing cost included in the cost of goods sold is matched with the revenues. Under variable costing, all of the fixed manufacturing cost is deducted in the period in which it is incurred, regardless of the amount of inventory change. Thus, when inventory decreases, the absorption costing income statement will have a lower income from operations.